Suppose the current market rate of interest is 22 percent. Ayşe is subject to a 40...

  • Suppose the current market rate of interest is 22 percent. Ayşe is subject to a 40 percent marginal tax rate on her interest income. What is Ayşe’s equilibrium marginal rate of time preference? Suppose the marginal tax rate John is subject to decreases to 30 percent. Can you predict the effect of the decrease in marginal tax rates on Ayşe’s current saving?
Answer #1

rate of time preference: it is the trade-off between present and future consumption. Value of present consumption is always weighed more than the value of future consumption. Hence forgoing consumption today would require getting something extra tomorrow. For eg., if Jim and Bob go out for a drink and Jim has no money so Bob lends Jim $10. The next day Bob comes back to Jim, and Jim says, "Bob, you can have $10 now, or at the end of the month when I get paid I will give you $15." Bob's time preference would change depending on if he trusted Jim and how much he needs the money now, thinks he can wait, or would prefer to have $15 at the end of the month than $10 now. Present and expected needs, present and expected income affect the time preference.

Equilibrium rate of time preference will be such that where the net return from savings equals the dis-utility from not consuming today. hence here, it will be = net return from savings = 0.22(Y) - 0.4*.022 (Y) that is rate of interest on income less the marginal tax on interest income. It will be, 022-0.088= .132
Hence equilibrium rate of time preference= 0.132 or 13.2%

With a decrease in tax rate, the net return on savings will go up. Hence her current savings will increase as she will now get even more in future by sacrificing one unit of her present consumption today. Her net return will now = 0.7*0.22= .154 or 15.4%